RECENT PROGRESS IN THE ROTATIONAL SPECTROSCOPY OF NITRIC ACID

\underline{DOUGLAS T. PETKIE}, Department of Physics, Wright State University, Dayton OH 45435; PAUL HELMINGER, Department of Physics, University of South Alabama, Mobile, AL 36688; SIDNEY H. YOUNG, Department of Chemistry, University of South Alabama, Mobile, AL 36688; R. A. H. BUT-LER and FRANK C. DE LUCIA, Department of Physics, The Ohio State University, Columbus, OH 43210.

Nitric acid has received spectroscopic attention for its internal rotation and its role in atmospheric remote sensing of the stratosphere. Current studies of the millimeter and submillimeter wave rotational spectrum of several vibrational states lying above 1000 cm^{−1} will be presented that are of theoretical and atmospheric interest. The breadth of assignments and the corresponding analyses vary with each state and in most cases include assignments made in the symmetric-top limit. The first group of states include ν_{8} + ν_{9}, ν_{6} + ν_{7}, ν_{6} + ν_{9}, and ν_{7} + ν_{9}, all of which exhibit a torsional splitting and have well developed analyses. Tentative assignments and analyses of several other states include 2ν_{7}, 3ν_{9}, ν_{3}, ν_{4} and possibly ν_{5} + ν_{9} will also be presented, several of which also exhibit a torsional splitting. For states in which there also has been infrared work, the relationship with these pure rotational studies will be considered. Additionally, the analysis of the ground state rotational spectrum of H^{15}NO_{3} measured in natural abundance will also be presented.